OpenStudy (anonymous):

I have four math courses this semester please help me: Plane Trigonometry: Give all the six functions if sec(theta) is 7/5. Plane Geometry: Find the Total Area of a cube with 6 cm edge. Statistic: Age of workers 57, 44,53,44,45,52,49,48,41,52,55,58,57,51,59,42,39,41,42,45,46,47,47,42,44 Determine the age of workers belong to the 6th decile and interpret. Modern Geometry: Given Rectangle FROG with FR=2x+y, RO=x+2, OG=x+12 and FG=y+6. Find the area and perimeter. Just want to know. I just came from my Pre-final exams. Questions i encounter.

OpenStudy (anonymous):

Plane Geometry: Area of one cube = 6*6 = 36 cm^2. There are 6 faces in a cube. The total area is 6*36 = 216 cm^2.

OpenStudy (anonymous):



OpenStudy (anonymous):

using the picture above: perimeter = sum of all sides....so sum it up. Area of a rectangle is \[A = length * width\]